# Prospects for detection of ultra high frequency gravitational waves from compact binary coalescenses with resonant cavities

### v2

@@ IV Event rate
$\times$ $\left(\frac{0.4K}{T_{\text{sys}}}\right)^{\frac{3}{2}}\left(\frac{m}{10^{-14}M_{\odot}}\right)^{\frac{427}{148}}\left(\frac{t_{\text{tot}}}{1\ \text{yr}}\right).$
The prefactors for the other setups are $5.6\times 10^{-49}$ for the 0.34 GHz configuration and $5.2\times 10^{-52}$ for the 11.47 GHz configuration. It should first be noticed that the mass appears with the exponent 2.89. The event rate increases very fast with higher masses. This was not a priori entirely obvious as 2 different effects do favour small masses. First, as previously explained, the signal spends a longer amount of time within the experimental bandwidth for light black holes. Second, the number density of black holes (and therefore of mergings) per unit time and volume increases – for a fixed total density – for small masses. Still, both those effects are far from compensating the stress increase with large masses.
In the case where the limit is the duration of the experiment (or, more exactly of the “run”) in the previous taxonomy, one is led to:
$N\approx 3.4\times 10^{-61}$$\times$ $\left(\frac{\nu}{2.67\ \text{GHz}}\right)^{\frac{23}{4}}\left(\frac{\eta}{0.1}\right)^{3}\left(\frac{B_{0}}{27\ \text{T}}\right)^{3}\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\text{m}^{3}}\right)^{\frac{5}{2}}\left(\frac{Q}{10^{5}}\right)^{\frac{9}{4}}$
$\times$ $\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{3}{2}}\left(\frac{m}{10^{-17}M_{\odot}}\right)^{\frac{153}{37}}\left(\frac{t_{\text{max}}}{60\ \text{s}}\right)\left(\frac{t_{\text{tot}}}{1\ \text{yr}}\right).$
The prefactors for the 0.34 GHz and 11.47 GHz setups are respectively equal to $1.7\times 10^{-61}$ and $1.8\times 10^{-61}$. As expected, the mass now enters the game with an even higher power of order 4.14. It should also be underlined that two different time scale are in order to estimate the event rate.
In the case where the limit is driven by the sampling rate, the number of events becomes
$N\approx 2.5\times 10^{-39}$$\times$ $\left(\frac{\nu}{2.67\ \text{GHz}}\right)^{\frac{5}{2}}\left(\frac{\eta}{0.1}\right)^{3}\left(\frac{B_{0}}{27\ \text{T}}\right)^{3}\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{2}}$
$\times$ $\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{3}{2}}\left(\frac{m}{10^{-9}M_{\odot}}\right)^{\frac{121}{74}}\left(\frac{t_{\text{tot}}}{1\ \text{yr}}\right).$
The prefactors are $1.0\times 10^{-36}$ for the 0.34 GHz configuration and $1.2\times 10^{-41}$ for the 11.47 GHz configurations.
The mass dependence is the weakest one, with an exponent close to 1.64. This is directly due to the fact that the positive power entering the maximum distance is smaller, because of the “saturation” effect induced by the limited sampling rate.
Finally, in the case of a phase-sensitive instrument such as a lock-in amplifier the number of event writes:
$N\approx 2.1\times 10^{-55}$$\times$ $\left(\frac{\nu}{2.67\ \text{GHz}}\right)^{\frac{5}{2}}\left(\frac{\eta}{0.1}\right)^{3}\left(\frac{B_{0}}{27\ \text{T}}\right)^{3}\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{2}}$
$\times$ $\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{3}{2}}\left(\frac{m}{10^{-17}M_{\odot}}\right)^{\frac{121}{74}}\left(\frac{t_{\text{tot}}}{1\ \text{yr}}\right).$
The prefactors are $8.6\times 10^{-53}$ for the 0.34 GHz configuration and $9.7\times 10^{-58}$ for the 11.47 GHz configurations.
The mass dependence is exactly the same than in the previous situation where the limit was driven by the sampling rate.
FIG. 6: Number of expected events per year with the GrAHal 2.67 GHz configuration as a function of the black holes mass, assuming that the upper limits on density is saturated. The black dashed line corresponds to an estimate with $t_{\text{max}}=1$ year whereas the purple dashed line corresponds to the lock-in case. The hatched region on the right is excluded by the results presented Fig. 1.
Figure 6 displays the number of expected events per year for the GrAHal 2.67 GHz configuration. This is a monotonically increasing function of the black hole mass. As expected from the estimations of the distances to mergers, all numerical values are extremely small. Analyses beyond this simple result are performed in the next section.
V Discussion
What conclusions can be drawn from the numbers previously obtained? Let us first focus on the event rate.
In the case of the GraHal experiment, for most of the parameter space, the first setup (low frequency range) is the best one and the last setup (high frequency range) is the worst one. This is not so surprising as the usual “figure of merit”, $B^{2}V$, often used to have a crude estimate of the performances, is indeed higher when going to bigger volumes (even though the magnetic field is decreased)  [^3] The current haloscopes constraints on the axion-photon coupling follow this trend, see https://cajohare.github.io/AxionLimits/docs/ap.html.. More importantly, one should notice that the way the experimental parameters enter the final estimate does depend on the considered regime – that is on the mass range. This is important to optimize the experimental design. Basically, we show that:
  • in all cases, the number of events scales as $B_{0}^{3}$, $V_{\text{cav}}^{2.5}$, and $T_{\text{sys}}^{-1.5}$,
  • for intermediate masses (signal drift limited case), the number of events also scales as $Q^{1.5}$ and $\nu^{3.75}$,
  • for small masses (duration of the experiment limited case), the number of events also scales as $Q^{2.25}$ and $\nu^{5.75}$,
  • for large masses (sampling rate limited case), the number of events scales as $\nu^{2.5}$ and does not depend on $Q$.
It might be welcome to increase the dimensionless coupling coefficient $\eta$, for example by considering the coupling with other resonant modes in the cavity, as studied in Berlin et al. (2022a). From the experimental viewpoint it might be the easiest way to increase the sensitivity. However, as $\eta$ is anyway bounded from above at 1, no dramatic enough improvement can be expected from this side alone.
In principle, increasing the sampling frequency might seem appealing as it would allow to extrapolate to the region $10^{-12}-10^{-6}M_{\odot}$ the (more favourable) slope of the curve obtained in the region $10^{-15}-10^{-12}M_{\odot}$, in Figs. 4, 5, and 6. This is however not possible beyond the value chosen here, which is optimum for a fixed spectral width $\Delta\nu$, that is at given quality factor (once $\nu$ is set). Interestingly, for large masses, there is no Q-dependence in the calculated sensitivity, distance, and event rates: what is won on the one hand is lost on the other hand. However, the maximal chirp mass $\mathcal{M}_{\text{max}}^{\Delta\nu}$ corresponding to the upper end of the “favourable” regime does depend on Q. When inserting the value of $t_{min}$ saturating the Shannon-Nyquist limit in Eq. (III.1.1), one sees that $\mathcal{M}_{\text{max}}^{\Delta\nu}\propto Q^{-6/5}$. This does not come as a surprise as this would increase $\Delta\nu$ and therefore allows for a higher sampling rate. Decreasing $Q$ does however reduce the sensitivity in the intermediate range.
Care should be taken with the $\nu$-dependence as it also changes the maximum (chirp) mass to which the experiment is sensitive. This variable can therefore not be used as naively as the others.
Beyond those scaling laws, the main lesson is that the actual numbers are tiny. We emphasize that no short-run improvement – at least technologically – should raise the event rate near to anything observable. This is one of the main conclusions of this work. Not to mention that several optimistic hypotheses were done. First, it was assumed that the abundance of black holes in the low-mass range correspond to the observed limits, around one percent of the critical density Niikura et al. (2019b). Moreover, it is also implicitly assumed that most black holes have masses such that the inspiralling signal falls within the (very narrow) experimental bandwidth.
Finally, it should also be reminded, as well known since decades, that the naive extrapolation of the primordial power spectrum observed at large scales, in the cosmological microwave background, leads to a density contrast at small scales which is, by far, too small to produce any black hole. Many smart ways out of this simple conclusion were investigated (see, e.g., Carr and Kuhnel (2020) and references therein), especially taking into account quantum diffusion in single-field models of inflation Ezquiaga et al. (2018), leading to a large non-Gaussian exponential tail that enhances the probability of collapse of primordial black holes even for small amplitudes of the power spectrum Ezquiaga et al. (2020). However, recent works tend to show that models producing appreciable amounts of primordial black holes generically induce too large one-loop corrections on large scales Kristiano and Yokoyama (2022). This claim has been questioned, particularly in the case of quantum diffusion, but it should anyway be kept in mind that even if the expected event rate was much higher, the way such black holes would be formed would still remain very speculative.
Another point of view could however be advocated: it would consist in not focusing on an “expected” rate but, instead, in investigating which mergers could be detected without violating existing observational constraints, independently of the underlying production mechanism and naively extrapolated cosmological density. By constraints, we mean here direct conflicts with data. The usual limits on the density of primordial black holes indeed rely on the assumption of a smooth distribution following a typical dark matter halo profile. They can be evaded if one assumes a local over-density in the form of PBH clusters, as follows from large non-Gaussian tails Ezquiaga et al. (2022). This is however possible and phenomenological potentially relevant on this sole basis. A (controversial) argument has even been given to suggest that than some trans-Neptunian anomalies and lensing events might be due to an unexpectedly high amount of local light black holes Scholtz and Unwin (2020).
However, even this optimistic scenario remains out of reach. Figure 5 indeed shows a kind of conspiracy. Whatever the considered chirp mass, the objects would have to be so close to the Earth (for this mass) that they should already have been detected, either directly or through their gravitational disturbances. Direct detection via e.g. X-ray emission from gas falling onto them is however not feasible, due to the extremely low flux received on Earth, even at solar system distances. Planet 9 is an example of gravitational disturbance. We leave a detailed study for a future work.
Clearly, the situation is less dramatic than in the purely consistent dark matter view presented before, and gaining a few orders of magnitude could allow the experiments to reach an interesting sensitivity. The most interesting part is the “high mass” one, that is the one corresponding to the sampling frequency-limited case. As previously shown, at fixed value of $\nu$, the maximum distance for detection scales as $B_{0}\times\eta\ \times V_{vac}^{0.83}\times T^{-0.5}$ and does not depend on $Q$. This underlines where efforts should be focused and their relative benefits. Of course, for evaluation the possible number of events, one should take the cube of the previous expression but this does not change the relative weights.
Note that we have assumed throughout that the detected gravitational waves come from a circularized orbit of coalescing compact binaries. However, for the small masses we are considering here, the circularization via emission of GW is very inefficient, so one should consider instead highly eccentric orbits Maggiore (2007). The power emitted by GW in eccentric orbits has an extra factor $F(e)=(1-e^{2})^{-7/2}(1+73/24\,e^{2}+37/96\,e^{4})$, so that a very eccentric orbit, with $e\simeq 1$, will emit much more energy per period and thus the connection between strain, distance, masses and power needs to be reevaluated. We leave for a future publication the analysis of such interesting cases. Very roughly speaking, to be able to efficiency detect those bursts, one would need a strong hierarchy between the orbital frequency of the system and the mean frequency of the emitted gravitational waves, which is not impossible. In this case, one could optimise the detector spectral width just in-between the physically relevant frequencies. This would both allow to have a time sampling of the orbit, therefore taking advantage of the temporal structure of the signal, and to select the resonance frequency at the expected signal frequency. If the signal is spread in Fourier space, much will inevitably be lost.
VI Other sources
@@
Beyond black holes, quite a lot of compact objects could, in principle, be considered. Among other sources, one can think to conventional compact objects, boson stars, fermion stars, dark matter stars, multi-components stars, or dark energy stars Giudice et al. (2016). We do not wish to discuss here which ones are compatible with the frequency range considered here and which ones are not. Neither do we investigate the naturalness of those sources. All the relevant models are anyway quite exotic and speculative. We therefore simply discuss briefly the effect of introducing a compactness parameter, in the sense of Alho et al. (2022), $C=Gm/(Rc^{2})$ – assumed to be the same for the two merging bodies, hence the unspecified mass $m$ and size $R$ for each object.
@@
For black holes, the compactness is $C=1/2$. It can only be smaller for other sources. This obviously makes the situation even worst. In general, the radius of the innermost stable circular orbit is
@@
where, as before, $m$ is the total mass of the considered binary system. As for black holes, the ISCO determines the end of the inspiral phase and the beginning of the merging phase, hence the last stage where the approach considered here basically holds.
It is quite usual in the literature (see, e.g. Aggarwal et al. (2021)) to express the strain for such a system (assuming that the two objects have the same mass $m$) as
@@
It seems to be proportional to the compactness $C$. However, in a sense, this way of writing the strain is over-simplistic. The C-factor actually enters this expression of the strain only because (at leading order) it is implicitly written at the ISCO, which is obviously reached earlier for a less compact system. In fact, Eq. (37) gives the maximum strain for a given inspiralling system. This is not the relevant quantity for the analysis presented here as, as explained in the previous sections, the narrow bandwidth of the experiments in the considered frequency range makes it highly unlikely that the system is observed at the most favourable stage of its evolution.
@@
where $m_{\text{max}}$ is the maximum allowed individual mass previously estimated. Within this limit (which is stringent for non-compact sources), the previous strain calculation, and therefore the distance estimates, still hold without correcting for the $C-$factor. One should keep in mind that for most non-exotic objects the ISCO is located inside the considered object!
VII Conclusion
The fact that haloscopes, designed to search for axions, might allow the detection of ultra high frequency gravitational waves is quite amazing. This opens an exciting new window on the Universe. This is obviously a promising and rapidly developing field with appealing features to search for new physics. This might lead to surprises in astrophysics, cosmology, or particle physics.
There might exist interesting stochastic signals, like backgrounds from the early Universe (see references in Aggarwal et al. (2021)), detectable by microwave cavities Herman et al. (2022). It could also be that close hyperbolic encounters García-Bellido and Nesseris (2017, 2018); Morrás et al. (2022) lead to measurable gravitational waves in the frequency band considered here. Several other exotic ideas also deserve consideration. We leave their detailed investigation (among many other possible sources) for future works. In particular, another possibility for detection with these resonant microwave cavities is the redshifted spectrum of GW coming from preheating García-Bellido and Figueroa (2007); García-Bellido et al. (2008). This spectrum has sufficient amplitude at high frequencies, where there are no expected astrophysical signals, that it is worth exploring.
@@ VII Conclusion
We have performed an in-depth estimate of the sensitivity in all physical regimes and pointed out where experimental improvements would be efficient. We have underlined strong correlations between parameters that are sometimes inappropriately considered as independent and have shown how to take them into account. In particular, temporal aspects have been carefully accounted for. We have calculated distances that can be probed and the related possible rates of events. Current experiments remain very far from the relevant sensitivity.
@@
It would certainly be welcome to investigate in more details the signal-to-noise ratio which has been estimated in an extremely crude way in this work. Although this would, quite certainly, not alter our conclusions, this might be important when dealing with more favourable physical situations. Obviously, a detailed time analysis of the signal would also be important. This should be considered in the future but this study was devoted to the investigation of the way axion haloscopes could be directly used for gravitational waves.
@@
To summarize, the main new inputs of this work – when compared with previous studies – are the following. First, the frequency of the experiment does not fix the mass of the binary systems that could, in principle, be probed: a double degeneracy is at play. Second, and more importantly, the duration of the physical signal in the bandwidth of the instrument can be extremely small and taking this into account drastically reduces the usually assumed sensitivity, hence the number of expected events. This is especially relevant for systems generating the higher strain. Third, for those systems, the limited sampling rate – associated with the fixed bandwidth – does not even allow to measure the signal optimally during this very brief time. This reduces even further the sensitivity.
VIII Acknowledgements

### v3

@@ III.2 Optimized resolution bandwidths
$\Leftrightarrow D$ $<$ $7.4\times 10^{4}\times\bigg{(}\frac{\nu}{2.67\ \text{GHz}}\bigg{)}^{\frac{5}{6}}\left(\frac{\eta}{0.1}\right)\left(\frac{B_{0}}{27\ \text{T}}\right)\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{6}}\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-14}M_{\odot}}\right)^{\frac{5}{6}}~{}\text{m}..$
IV Results and discussion
Figure 4 shows the strain sensitivity as a function of the chirp mass in all the considered regimes. The plain black line corresponds to the most “realistic” case whereas dashed lines correspond to the different hypotheses discussed previously. It should be first emphasized once more that the strain required for the considered binary system to be detected does depend upon the (chirp) mass. This is because the time structure of the signal – hence its interplay with the different timescales of the experiment – highly depends on the mass. It is therefore not sufficient to characterize a cavity experiment by a (minimum) strain to which it is sensitive. As we show here, for masses between $10^{-18}$ and $10^{-6}$ solar mass, the strain required for the signal to be detectable varies by 10 orders of magnitude. The higher the mass, the higher the required strain (as the signal becomes shorter). Impressively, for the highest detectable masses, corresponding to a system seen very close to merging, the required strain for the measurement to be possible is very high, around $h\sim 10^{-11}$, and very different from the sometime advertised sensitivity. The situation is obviously not the same for broadband experiment like Virgo and LIGO (see, e.g., Section 7 of Maggiore (2007)).
FIG. 4: Strain sensitivity of the GrAHal 2.67 GHz configuration as a function of the chirp mass of the physical system. The three different temporal regimes can be easily identified. The region above $10^{-6}$ solar mass is forbidden. The dashed line on the left part corresponds to an improvement relying on an optimized resolution bandwidth. The three dashed lines on the right part correspond to the different temporal effects – accounted for individually – described in the text. The dark plain line corresponds to the best possible estimate taking everything into account. The values $\Delta\nu^{-1}=10^{-4}$ s and $t_{\text{max}}=60$ s have been chosen.
Figure 5 displays the maximum distance at which the binary system can de detected as a function of the chirp mass. The plain line still corresponds to the “most” realistic case. Although the required strain is higher for larger masses, the distance remains an increasing function of the mass because the fast varying amplitude of signal plays a bigger role than the temporal effects, shown on the previous figure, which favors small masses. This was not a priori obvious. Interestingly, one can however notice that the accessible distance remains constant above a few times $10^{-12}$ solar mass. This is because, in this regime, the subtle experimental effects happen to exactly compensate for the physical increase in generated strain. We have previously argued that, should the sensitivity be smaller for smaller masses, it would still be mandatory to consider them as we do not know what are the actual masses of hypothetical light black holes around us. But this figure gives another argument: the sensitivity (in distance) happens to be flat between $2\times 10^{-12}$ and $10^{-6}$ solar masses.
FIG. 5: Maximum distance to the inspiralling system for the GrAHal 2.67 GHz configuration as a function of the chirp mass of the physical system. The three different temporal regimes can easily be identified. The region above $10^{-6}$ solar mass is forbidden. The dashed line on the left part corresponds to an improvement relying to an optimized resolution bandwidth. The three dashed lines on the right part correspond to the different temporal effects – accounted for individually – described in the text. The dark plain line corresponds to the best possible estimate taking everything into account. The values $\Delta\nu^{-1}=10^{-4}$ s and $t_{\text{max}}=60$ s have been chosen.
The most important and obvious result from this study is clear: beyond interesting “trends” revealing nice physical effects, the numerical value of the maximum distance is tiny. At best, it reaches 1000 km. We are forced to conclude that no detection can be expected. It is not even worth calculating an expected event rate or trying to identify which parameters should be improved.
If one however plays the game to follow Clesse and García-Bellido (2017) and consider primordial black holes clustered in virialized sub-halos in a model fully characterized by the local density contrast $\delta$ (that can be taken close to $10^{9}$ for a rough order of magnitude), one is led to an event rate $\tau\sim 10^{-56}~{}\text{yr}^{-1}$. Even if the formula we have used in this work for the SNR is deeply sub-optimal it seems hard to imagine that such a tiny number can be compensated for by any refined analysis.
V Other sources
@@
Beyond black holes and ignoring the extremely small numerical values obtained previously, quite a lot of compact objects could, in principle, be considered. Among other sources, one can think to conventional compact objects, boson stars, fermion stars, dark matter stars, multi-components stars, or dark energy stars Giudice et al. (2016). We simply discuss here briefly the consequences of introducing a compactness parameter, in the sense of Alho et al. (2022), $C=Gm/(Rc^{2})$ – assumed to be the same for the two merging bodies, hence the mass $m$ and size $R$ – for purely theoretical purposes, in case this study is used in another context.
@@
For black holes, the compactness is $C=1/2$. It can only be smaller for other sources. This obviously makes the situation even worse. In general, the radius of the innermost stable circular orbit is
@@
where, as before, $m$ is the total mass of the considered binary system. It is quite usual in the literature (see, e.g. Aggarwal et al. (2021)) to express the strain for such a system (assuming that the two objects have the same mass $m$) as
@@
It seems to be proportional to the compactness $C$. However, in a sense, this way of writing the strain is over-simplistic. The C-factor actually enters this expression of the strain only because (at leading order) it is implicitly written at the ISCO, which is obviously reached earlier for a less compact system. In fact, Eq. (31) gives the maximum strain for a given inspiralling system. This is not the relevant quantity for the analysis presented here as, as explained in the previous sections, the narrow bandwidth of the experiments in the considered frequency range makes it highly unlikely that the system is observed at the most favorable stage of its evolution and as, in addition, the full sensitivity does not even make this situation easier to detect.
@@
where $m_{\text{max}}$ is the maximum allowed individual mass previously estimated. Within this limit (which is stringent for non-compact sources), the previous strain calculation, and therefore the distance estimates, still hold without correcting for the $C-$factor. One should keep in mind that for most non-exotic objects the ISCO is located inside the considered object.
VI Conclusion
The fact that haloscopes, designed to search for axions, might allow the detection of ultra high frequency gravitational waves is quite amazing. This opens an exciting new window on the Universe. This is obviously a rapidly developing field with appealing features to search for new physics which might lead to surprises in astrophysics, cosmology, or particle physics.
There might exist interesting stochastic signals, like backgrounds from the early Universe (see references in Aggarwal et al. (2021)), detectable by microwave cavities Herman et al. (2022). Signals from the late Universe can also be expected Aggarwal et al. (2021). Among others Arvanitaki et al. (2015); Brito et al. (2015); Tootle et al. (2022); Casalderrey-Solana et al. (2022), it could also have been hoped that close hyperbolic encounters García-Bellido and Nesseris (2017, 2018); Morrás et al. (2022) lead to measurable gravitational waves in the frequency band considered here. This seems not to be the case Teuscher et al. (2024); Barrau et al. (2024). Several other exotic ideas also deserve consideration. We leave their detailed investigation (among many other possible sources) for future works.
@@ VI Conclusion
We have performed an in-depth estimate of the sensitivity in all physical regimes and underlined strong correlations between parameters that are sometimes inappropriately considered as independent and have shown how to take them into account. In particular, temporal aspects have been carefully accounted for. We have calculated distances that can be probed and ours results exhibit that current experiments remain very far from the relevant sensitivities. This is our main conclusion: at this stage, no signal at all is to be expected from binary systems of black holes.
@@
It would certainly be welcome to investigate in more details the signal-to-noise ratio which has been estimated with the Dicke formula in this work. Although this would, quite certainly, not alter our conclusions, this might be important when dealing with more favorable physical situations. Obviously, a detailed time analysis of the signal would also be important and might slightly change the picture. This should be considered in the future. However, once again, this study was devoted to the investigation of the way axion haloscopes could be directly reused to search for gravitational waves.
In addition, it is worth pointing out that our analysis is stricto sensu valid only for masses quite below the upper bound, that is for systems seen long before the merging. Not only because post-Newtonian corrections are ignored but also because, even at the Newtonian level, it is not correct to assume, as we did, that the orbit remains circular – albeit with a time-dependent radius – close to the merging. The fact is that the trajectory is not a circle anymore just before the coalescence: there is no circle even approximately tangent to the trajectory. We do not expect any correction from the detailed shape of the orbit to drastically change our conclusions (the Fourier transform of the signal should even be wider, making the situation even worse) but, in principle, the present analysis is fully reliable only when $\mathcal{M}_{C}\ll\mathcal{M}_{C}^{max}$ (the latter corresponding to a gravitational wave frequency at merging equal to the cavity resonant one).
@@
To summarize, the main new inputs of this work – when compared with previous studies – are the following. First, the frequency of the experiment does not fix the mass of the binary systems that could, in principle, be probed: a double degeneracy is at play. Second, and more importantly, the duration of the physical signal in the bandwidth of the instrument can be extremely small and taking this into account drastically reduces some of the usually assumed sensitivities, hence the number of expected events. This is especially relevant for systems generating the higher strain. Third, the combined effect of the noise averaging associated with the limited sampling rate with the partial cavity charging leads to a decrease of the effective quality factor reducing further – by many orders of magnitude – the sensitivity. Interestingly, when all the experimental effects are taken into account, the maximum probed distance does not depend anymore on the mass above $10^{-12}$ solar mass.
The numerical values we have obtained seem to close the window on compact binary systems as sources of GHz gravitational waves detectable by resonant cavities.
Beyond this somehow disapointing conclusion, we have tried to make this study as pedagogical as possible so that the methodology can be easily re-used for other setups.
VII Acknowledgements